# Ordering the topological order in the fractional quantum Hall effect

### v2

@@ C.5 Examples
    When the theory ${\cal T}$ can be factorized as ${\cal M}\boxtimes\{1,c\}$ where ${\cal M}$ is an MTC, a minimal modular extension is ${\cal M}\boxtimes\{1,c,\bm{m},\bm{m}c\}$. Then there exists an Abelian $b\in{\cal M}$ such that $\bm{\gamma}=\bm{m}\times b$, and $e^{i\pi Q(a)}=e^{2\pi iB(\bm{m}\times b,a)}=e^{2\pi iB(b,a)}$ for $a\in{\cal M}$, since $B(\bm{m},a)=0$. In other words, $Q(a)=2B(b,a)$ mod 2. Note that the vison $v=\bm{\gamma}^{2}=\bm{m}^{2}b^{2}$. Depending on the minimal modular extension, $\bm{m}^{2}$ may be $1$ or $c$.
@@
    In particular, if $v=1$, then we must have $b^{2}=1$, i.e., $b$ generates a $\mathbb{Z}_{2}^{(1)}$ one-form symmetry. In this case, (C.34) can be viewed as a mapping from $\cal T$ to another bosonic TQFT, which is also the one discussed in Cor. C.18. A similar map between bosonic TQFTs was studied in [167] as Lorentz symmetry fractionalization (because attaching a transparent fermion changes the spin by $1/2$).
@@
    For a non-split example, consider the Moore-Read state $\mathsf{Pf}_{q,1,1}^{f}$ (the physical realization has $\sigma_{H}=\frac{1}{q}$ and $c_{-}=\frac{3}{2}$). $\psi_{2q}$ is the transparent fermion. The unique minimal modular extension with $c_{-}=\frac{3}{2}$ is $\mathcal{T}^{\rm ext}={\rm Ising}\boxtimes\mathcal{A}^{4q,1}$. In this theory The Abelian $\pi$ flux anyons $\mathcal{A}_{1}$ are $1_{n}$ and $\psi_{n}$ with $n=1,3,\dots,4q-1$, where we use the notation of App. A.4. One can take $\mathcal{A}_{1}/\{1,c\}=\{1_{n}|n=1,3,\dots,4q-1\}$. Suppose $\bm{\gamma}=1_{n}$ (so the vison $v=\bm{\gamma}^{2}=1_{2n}$). Using the relation $h(\bm{\gamma})=\frac{\sigma_{H}}{8}$ mod 1, we find $\sigma_{H}=\frac{n^{2}}{q}$ mod 8. The non-Abelian anyon $\sigma_{1}$ then has charge $Q(\sigma_{1})=\frac{n}{2q}$ mod 2.
@@ C.6 Restriction to symmetry lines
First, it has its own vison $\bm{\gamma}$ generating ${\cal V}^{s^{\prime},r^{\prime}}$ with some integers $s^{\prime}$ and $r^{\prime}$ to be determined. Also, $c$ generates ${\cal V}^{2,2}$. Comparing with the discussion in App. b, the anyons in ${\cal T}^{\rm ext}_{0}$ are even under the $\mathbb{Z}_{2}^{(1)}$ one-form symmetry generated by $c$ and the anyons in ${\cal T}^{\rm ext}_{1}$ are odd under that symmetry.
@@
$\begin{split}&h(v^{j})={rj^{2}\over 2s}\text{ mod }{1\over 2}\\ &Q(v^{j})=\left(2Q^{\rm ext}(v^{j})\right)\text{ mod }1={rj\over s}\text{ mod }1\\ &\hat{h}(v^{j})=\left(h(v^{j})+{1\over 2}Q(v^{j})\right)\text{ mod }1=\left({rj^{2}\over 2s}+{rj\over 2s}\right)\text{ mod }1\\ &r,s\in 2\mathbb{Z}+1\,.\end{split}$ (C.37)
@@
$\begin{split}&h(v^{j})={rj^{2}\over 2s}\text{ mod }{1\over 2}\\ &Q(v^{j})=\left(2Q^{\rm ext}(v^{j})\right)\text{ mod }1={rj\over s}\text{ mod }1\\ &\hat{h}(v^{j})=\left(h(v^{j})+{1\over 2}Q(v^{j})\right)\text{ mod }1=\left({rj^{2}\over 2s}+{rj\over 2s}\right)\text{ mod }1\\ &r\in 2\mathbb{Z}\,.\end{split}$ (C.38)
@@ Appendix D Gauging $\mathbb{Z}_{\ell}^{(1)}$ reduces the total quantum dimension by $\ell$
Before going to the proof, we need a lemma, which is of independent interest and is used in other places in the text:
@@
[57] Suppose $\cal T$ has $G$-graded fusion rules where $G$ is a finite group, i.e., ${\cal T}=\oplus_{g\in G}{\cal T}_{g}$ and ${\cal T}_{g}\times{\cal T}_{h}\in{\cal T}_{gh}$, then ${\cal D}_{g}^{2}=\sum_{a_{g}\in{\cal T}_{g}}d_{a_{g}}^{2}={\cal T}_{0}^{2}$, where 00 denotes the identity element of the group $G$. In particular, ${\cal D}_{g}$ is independent of $g$.
@@
We can also derive this result from the perspective of the partition function of the 2+1d TQFT. The three-sphere partition function before gauging is given by (8.3) ${\cal Z}_{\cal T}(S^{3})=S_{11}=\frac{1}{{\cal D}_{\cal T}}$. The partition function of the theory after gauging the $\mathbb{Z}_{\ell}^{(1)}$ 1-form symmetry on a closed manifold ${M^{(3)}}$ is given by (3.27)
@@
Finally, we can also prove Theorem D.1, using the RCFT perspective. We are interested in gauging the one-form symmetry $\mathbb{Z}_{\ell}^{(1)}$ generated by $b$ to map the TQFT $\cal T$ to ${\cal T^{\prime}}={\cal T}/\mathbb{Z}_{\ell}^{(1)}$. Since this one-form symmetry is anomaly-free, its elements $b^{j}$ braid trivially with each other, and therefore in terms of the elements of the $S$-matrix, $S_{b^{j},b^{i}}=S_{11}$.
@@ Appendix E Minimal modular extensions of $\mathcal{A}^{q\ell,p\ell}$
In this appendix, we construct an explicit Abelian minimal modular extension of $\mathcal{A}^{q\ell,p\ell}$, which we call ${\cal W}^{q,p,\ell}$. Other minimal modular extensions of $\mathcal{A}^{q\ell,p\ell}$ can be constructed using ${\cal W}^{q,p,\ell}$ and equation (8.13).
For simplicity, we limit ourselves to bosonic theories. Then, $pq\ell$ has to be even, and there are no further restrictions due to the global $\mathrm{U}(1)$ symmetry.
@@
${\cal W}^{q,p,\ell}=\left\{\begin{array}[]{cc}\mathcal{A}^{q\ell_{1}^{2},p\ell_{2}^{2}}\boxtimes D(\mathbb{Z}_{\ell_{2}}),&pq\in 2\mathbb{Z},\\ \mathcal{A}^{q\ell_{1}^{2},p\ell_{2}^{2}}\boxtimes\mathcal{A}^{\ell_{2}^{2},1},&pq\in 2\mathbb{Z}+1\,.\end{array}\right.$ (E.2)

### v3

@@ C.5 Examples
[…]
@@
    In particular, if $v=1$, then we must have $b^{2}=1$, i.e., $b$ generates a $\mathbb{Z}_{2}^{(1)}$ one-form symmetry. In this case, (C.34) can be viewed as a mapping from $\cal T$ to another bosonic TQFT, which is also the one discussed in Corollary C.18. A similar map between bosonic TQFTs was studied in [165] as Lorentz symmetry fractionalization (because attaching a transparent fermion changes the spin by $1/2$).
@@
    For a nonsplit example, consider the Moore-Read state $\mathsf{Pf}_{q,1,1}^{f}$ (the physical realization has $\sigma_{H}=\frac{1}{q}$ and $c_{-}=\frac{3}{2}$). $\psi_{2q}$ is the transparent fermion. The unique minimal modular extension with $c_{-}=\frac{3}{2}$ is $\mathcal{T}^{\rm ext}={\rm Ising}\boxtimes\mathcal{A}^{4q,1}$. In this theory the Abelian $\pi$ flux anyons $\mathcal{A}_{1}$ are $1_{n}$ and $\psi_{n}$ with $n=1,3,\dots,4q-1$, where we use the notation of Appendix A.4. One can take $\mathcal{A}_{1}/\{1,c\}=\{1_{n}|n=1,3,\dots,4q-1\}$. Suppose $\bm{\gamma}=1_{n}$ (so the vison $v=\bm{\gamma}^{2}=1_{2n}$). Using the relation $h(\bm{\gamma})=\frac{\sigma_{H}}{8}$ mod 1, we find $\sigma_{H}=\frac{n^{2}}{q}$ mod 8. The non-Abelian anyon $\sigma_{1}$ then has charge $Q(\sigma_{1})=\frac{n}{2q}$ mod 2.
@@ C.6 Restriction to symmetry lines
First, it has its own vison $\bm{\gamma}$ generating ${\cal V}^{s^{\prime},r^{\prime}}$ with some integers $s^{\prime}$ and $r^{\prime}$ to be determined. Also, $c$ generates ${\cal V}^{2,2}$. Comparing with the discussion in Appendix b, the anyons in ${\cal T}^{\rm ext}_{0}$ are even under the $\mathbb{Z}_{2}^{(1)}$ one-form symmetry generated by $c$ and the anyons in ${\cal T}^{\rm ext}_{1}$ are odd under that symmetry.
@@
[…]
@@
[…]
@@ Appendix D Gauging $\mathbb{Z}_{\ell}^{(1)}$ reduces the total quantum dimension by $\ell$
Before going to the proof, we need a lemma from Ref. [33], which is of independent interest and is used in other places in the text:
@@
Suppose $\cal T$ has $G$-graded fusion rules where $G$ is a finite group, i.e., ${\cal T}=\oplus_{g\in G}{\cal T}_{g}$ and ${\cal T}_{g}\times{\cal T}_{h}\in{\cal T}_{gh}$, then ${\cal D}_{g}^{2}=\sum_{a_{g}\in{\cal T}_{g}}d_{a_{g}}^{2}={\cal T}_{0}^{2}$, where 0 denotes the identity element of the group $G$. In particular, ${\cal D}_{g}$ is independent of $g$.
@@
We can also derive this result from the perspective of the partition function of the 2+1d TQFT. The three-sphere partition function before gauging is given by (8.3) ${\cal Z}_{\cal T}(S^{3})=S_{11}=\frac{1}{{\cal D}_{\cal T}}$. The partition function of the theory after gauging the $\mathbb{Z}_{\ell}^{(1)}$ one-form symmetry on a closed manifold $M^{(3)}$} is given by (3.27)
@@
Finally, we can also prove Theorem D.1, using the RCFT perspective. We are interested in gauging the one-form symmetry $\mathbb{Z}_{\ell}^{(1)}$ generated by $b$ to map the TQFT $\cal T$ to ${\cal T^{\prime}}={\cal T}/\mathbb{Z}_{\ell}^{(1)}$. Since this one-form symmetry is anomaly free, its elements $b^{j}$ braid trivially with each other, and therefore in terms of the elements of the $S$-matrix, $S_{b^{j},b^{i}}=S_{11}$.
@@ Appendix E Minimal modular extensions of $\mathcal{A}^{q\ell,p\ell}$
In this appendix, we construct an explicit Abelian minimal modular extension of $\mathcal{A}^{q\ell,p\ell}$, which we call ${\cal W}^{q,p,\ell}$. Other minimal modular extensions of $\mathcal{A}^{q\ell,p\ell}$ can be constructed using ${\cal W}^{q,p,\ell}$ and Eq. (8.13).
For simplicity, we limit ourselves to bosonic theories. Then, $pq\ell$ has to be even, and there are no further restrictions caused by the global $\mathrm{U}(1)$ symmetry.
@@
${\cal W}^{q,p,\ell}=\left\{\begin{array}[]{cc}\mathcal{A}^{q\ell_{1}^{2},p\ell_{2}^{2}}\boxtimes D(\mathbb{Z}_{\ell_{2}}),\qquad&pq\in 2\mathbb{Z},\\ \mathcal{A}^{q\ell_{1}^{2},p\ell_{2}^{2}}\boxtimes\mathcal{A}^{\ell_{2}^{2},1},\qquad&pq\in 2\mathbb{Z}+1\,.\end{array}\right$. (E.2)